Opened 7 years ago

Closed 5 years ago

Last modified 8 months ago

#435 closed flight processing (fixed)

GB12_06, flight day 083a/2012, Harnhill

Reported by: knpa Owned by: knpa
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: adbe, knpa Other processors:

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/GB12_06-2012_083a_Harnhill

Data arrived from ARSF via network transfer on 26/03/2012

GB12_06 Scientific objective:

GB12_06 Priority: 8

PI: Anthony Beck

Sensors:

Camera (26/04/2012)
Eagle (07/06/2012)
Hawk (07/06/2012)
LiDAR (07/06/2012)

Change History (34)

comment:1 Changed 7 years ago by knpa

Hawk crashed on line 6 so there is no data for this line.

comment:2 Changed 7 years ago by knpa

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 low

comment:3 Changed 7 years ago by adbe

Started navigation processing

comment:4 Changed 7 years ago by adbe

Navigation processed

comment:5 Changed 7 years ago by knpa

  • Description modified (diff)
  • Priority changed from alpha 4 low to alpha 4 medium

comment:6 Changed 7 years ago by adbe

Processed RCD

comment:7 Changed 7 years ago by adbe

RCD ready for delivery check.

comment:8 Changed 7 years ago by knpa

  • Description modified (diff)

comment:9 Changed 7 years ago by adbe

Processing hyperspectral. Got the sct's:

line eagle hawk
1 -0.02 -0.06
2 -0.02 -0.06
3 -0.02 -0.06
4 -0.02 -0.06
5 -0.02 -0.06
6 -0.02 -----
7 -0.02 -0.06
8 -0.02 -0.06
9 -0.02 -0.06

comment:10 Changed 7 years ago by knpa

Starting RCD delivery check.

comment:11 Changed 7 years ago by knpa

No problems found.

Ready to deliver.

comment:12 Changed 7 years ago by adbe

  • Description modified (diff)

Camera delivered to Anthony Beck on 26/04/2012 on disk 108.

comment:13 Changed 7 years ago by adbe

Hyperspectral ready for delivery check

comment:14 Changed 7 years ago by knpa

Taking HS delivery check.

comment:15 Changed 7 years ago by knpa

We've not decided conclusively what we're doing with mapped BILs yet. May need to compress/provide a line splitter tool. Will DC the rest of the project in the meantime.

comment:16 Changed 7 years ago by knpa

All looks good apart from the following:

  1. Need to unix2dos the new view vector text files
  2. Logsheet has incorrect name format
  3. Should mention remaining SCT issue for the wobbly line in ReadMe
  4. ReadMe contents - flightlines subdirs not in alphabetical order
  5. Rename mapped files to correct name format

Will need to check again after you have zipped up the mapped files, and edited the ReadMe.

comment:17 Changed 7 years ago by adbe

  • Did unix2dos on text files
  • Renamed logsheet
  • Mentioned in ReadMe about wobbly line
  • Contents in alphabetical order
  • Renamed mapped files

comment:18 Changed 7 years ago by adbe

Processed LiDAR, LiDAR ready for delivery check.

comment:19 Changed 7 years ago by adbe

  • Owner set to adbe
  • Status changed from new to accepted

comment:20 Changed 7 years ago by knpa

Checked the things in the HS delivery that been updated - ReadMe, screenshots, zipped files.

Looks fine - ready to deliver.

comment:21 Changed 7 years ago by knpa

Taking LiDAR delivery check.

comment:22 Changed 7 years ago by knpa

Checked LiDAR. All looks fine apart from:

  • Incorrect file name format for ascii and las flightline files

comment:23 Changed 7 years ago by adbe

Renamed LiDAR and altered ReadMe.

comment:24 Changed 7 years ago by adbe

  • Cc adbe knpa added

comment:25 Changed 7 years ago by adbe

  • Description modified (diff)

Eagle, Hawk and LiDAR delivered to Anthony Beck on 07/06/2012 on disk 119.

comment:26 Changed 7 years ago by emca

There is full waveform data with this dataset that was not processed first time around. Currently processing this.

comment:27 Changed 7 years ago by emca

Full waveform lidar data sent to David Stott, Univ. of Leeds 25/06/2012

comment:28 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from adbe to knpa
  • Status changed from accepted to assigned

Beginning archiving.

comment:29 Changed 6 years ago by knpa

Made available to NEODC.

comment:30 Changed 6 years ago by knpa

Archived by NEODC. Moved to archive disk.

comment:31 Changed 5 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

comment:32 Changed 3 years ago by dac

Archiving

  • Moved deliveries to delivery. Hardlinked from discrete delivery to make final FW delivery complete.
  • Tidied up raw data and archiving to NEODC (currently they don't have a copy).
  • NEODC should also replace their lidar delivery (which is the two deliveries merged) with the final, complete delivery.
  • Uploading to NEODC.

comment:33 Changed 16 months ago by asm

JSON files - Lidar delivery

Problem with bng dem projection missing, solved using assign_projection.py -p '+proj=tmerc +lat_0=49 +lon_0=-2 +k=0.9996012717 +x_0=400000 +y_0=-100000 +ellps=airy +units=m +no_defs ' the_dem-bng.dem

comment:34 Changed 8 months ago by dac

Archiving

Full waveform data is available from NEODC.

Note: See TracTickets for help on using tickets.